# A Survey of Protocol Fuzzing

### v2

@@ 3.1.2. Constrained Testing Environment
Protocol fuzzing usually faces a constrained testing environment due to the tight coupling between protocols and the hardware. Firstly, numerous protocols are designed either for communications between low-level physical devices or for communications happened in specialized sectors, such as the protocols reside in the lower layers of the OSI refernce model, i.e., the physical and data link layers (Chen et al., 2005; 802, 2021, 2020; Alliance, 2015; SIG, 2016), or protocols designed for specific sectors like automotive (Zuo et al., 2021; AUTOSAR, 2016; OMG, 2018), industial control system (ICS) (Zhao et al., 2019; Bytes et al., 2022), eletricity grids, and aviation systems. In these cases, the testing throughput will be limited by the hardware dependencies, such as the lack of auotmation (Park et al., 2022; Song et al., 2022), the bottleneck for scalable fuzzing (Bytes et al., 2022; Ruge et al., 2020), etc. Besides, these physical dependencies also limit the application of advanced fuzzing techniques. This is because many advanced fuzzing techniques require greybox or whitebox testing information from the test target, which cannot be satisfied due to the lack of program analysis frameworks on these specific hardware (Zuo et al., 2022; Schepers et al., 2021; Zuo et al., 2021).
@@ 3.2. Summary of Existing Protocol Fuzzers
We have analyzed current protocol fuzzing research and encapsulated these efforts into a technical framework for protocol fuzzers, as depicted in Fig. 4. Note that the existing fuzzing works still follow the high-level concepts of general fuzzing but propose specific enhancements in these subcomponents to solve protocol-specific challenges. We first discuss the general concepts and functionalities of these components in this section, and then detail the specific enhancements existing works made in following sections (Sections 4 to 6).
A general fuzzer consists three basic components, namely input generator, executor and bug collector. In one iteration of fuzzing, input generator first produces a testing input to executor. Then executor executes the PUT with the given input and collects runtime information for the other two components. Finally the bug collector checks the runtime information to determine whether the input has triggered a bug.
Input generator. Ideally, this component is responsible for generating inputs to expose vulnerabilities inside PUTs as effective as possible. To realize this, protocol fuzzers usually implement input generator with three main subphases including communication model construction, scheduling and testcase construction. Communication model construction phase is responsible for learning the semantic constraints of the protocol to provide knowledge for the other phases. With protocol domain specific knowledge, scheduling phase decides all the scheduling configurations used in the next iteration of input generation, with an intention to expose more bugs. Testcase construction phase is responsible for producing testcase according to the scheduler’s instructions.
@@
Bug collector. The primary objective of the bug collector component is to enhance both the variety of detected bug types and the accuracy of these detections. The component is finely tuned to meticulously identify a broad spectrum of vulnerabilities, ranging from memory-safety bugs like buffer overflows to more subtle non-memory-safety bugs such as logic errors and specification violation.
@@ 4. Input Generator
In this section, we will introduce in detail how the existing works improve the input generator to solving the unique challenges in protocol fuzzing. As shown in Table 3, we summarize the techniques used by existing works in designing the three key phases in input generator. The covered works are selected from our paper set as long as their main techniques are directly related to the input generator. Besides the statistics of the mentioned three phases for these works, the table also lists their feedback information. According to Fig. 4, the feedback information can be provided by the executor or the bug collector. We only discuss how feedback information is used in input generator but leave the feedback collection related details for Section 5.2.
| Years | Work | T | Target | CMC | Scheduling |
Feedback | ||
2013 |
BED(Sesterhenn and Muench, 2013) | ● | General | Manual |
Sequential |
P |
State |
2013 |
Tsankov et al. (Tsankov et al., 2013) | ● | General | Manual |
P & S |
2014 |
Peach(Eddington, 2014) | ● | General | Manual |
Sequential |
P |
State |
2015 |
Pulsar(Gascon et al., 2015) | ● | General | TAPL |
SCHS |
P |
State |
2015 |
Beurdouche et al.(Beurdouche et al., 2015) | ● | TLS | Manual |
Random |
S |
State |
2015 |
Ruiter et al.(de Ruiter and Poll, 2015) | ● | TLS | TAAL |
Random |
S |
State |
2016 |
TLS-Attacker(Somorovsky, 2016) | ● | TLS | Manual |
Random |
P & S |
2016 |
Driller(Stephens et al., 2016) | ○ | General | - |
SPMS |
P |
Code Cov |
2017 |
Fan et al.(Fan and Chang, 2018) | ● | General | TAPL |
P & S |
2017 |
WiFuzz(Vanhoef, 2017) | ● | Wi-Fi | Manual |
Sequential |
P & S |
2018 |
TCPWN(Jero et al., 2018) | ● | TCP | Manual |
Sequential |
P & S |
State |
2018 |
IoTFuzzer(Chen et al., 2018) | ● | IoT (OASIS, 2019) | - |
P |
2018 |
Danial et al.(Daniel et al., 2018) | ● | OpenVPN | Manual |
Random |
S |
2018 |
DELTA(Lee et al., 2018) | ● | OpenFlow | Manual |
P & S |
2019 |
SeqFuzzer(Zhao et al., 2019) | ● | ICS | TAPL |
P |
State |
2019 |
Polar(Luo et al., 2019) | ◐ | ICS | - |
SPMS |
P |
Code Cov |
2019 |
IoTHunter (Yu et al., 2019) | ◐ | IoT | TAAL |
Sequential |
P |
Code Cov |
2019 |
MoSSOT (Shi et al., 2019) | ● | SSO(Group, 2018) | Manual |
Sequential |
P & GUI Ops |
2019 |
Chen et al. (Chen et al., 2019) | ◐ | General | - |
SPHS |
P |
State & Code Cov |
2019 |
Fuzzowski(Romero and Rivas, 2019) | ◐ | General | Manual |
P |
State & Code Cov |
2020 |
Walz et al.(Walz and Sikora, 2017) | ● | TLS | - |
Random |
P |
2020 |
DTLS-Fuzzer(Fiterau-Brostean et al., 2020, 2022) | ● | DTLS | TAAL |
Random |
S |
State |
2020 |
AFLNET (Pham et al., 2020) | ◐ | General | TAAL |
SRHS |
P |
State & Code Cov |
2020 |
SweynTooth (Garbelini et al., 2020) | ● | BLE(SIG, 2016) | Manual |
SPHS |
P & S |
State & #Bugs |
2020 |
Frankenstein (Ruge et al., 2020) | ◐ | Bluetooth | Manual |
Random |
P & S |
State & Code Cov |
2020 |
Peach* (Luo et al., 2020) | ◐ | ICS | - |
P |
Code Cov |
2020 |
aBBRate (Peterson et al., 2020) | ● | TCP | Manual |
Sequential |
S |
State |
2020 |
IJON (Aschermann et al., 2020) | ○ | General | PAL |
Random |
P |
State & Code Cov |
2020 |
FuSeBMC (Alshmrany and Cordeiro, 2020) | ○ | General | - |
Sequential |
P |
Code Cov |
2020 |
DPIFuzz (Reen and Rossow, 2020) | ● | QUIC(Iyengar and Thomson, 2021) | Manual |
Random |
P & S |
2020 |
Zou et al.(Zou et al., 2020) | ● | General | - |
P |
2021 |
ICSFuzzer (Fang et al., 2021) | ● | ICS (Corporation, 2020; Electric, 2009) | PAL |
SCHS |
P & GUI Ops |
State |
2021 |
StateAFL (Natella, 2022) | ◐ | General | PAL |
SPHS & SRHS |
P |
State & Code Cov & # of Bugs |
2021 |
TCP-Fuzz (Zou et al., 2021) | ◐ | TCP | Manual |
Random |
P & S & Syscall |
State Transition |
2021 |
Snipuzz(Feng et al., 2021) | ● | IoT | - |
P |
2021 |
Z-Fuzzer(Ren et al., 2021) | ◐ | Zigbee | - |
P & Interrupt |
Code Cov |
2021 |
PAVFuzz(Zuo et al., 2021) | ◐ | AV (AUTOSAR, 2016) | Manual |
Sequential |
P |
Code Cov |
2021 |
Aichernig et al.(Aichernig et al., 2021) | ● | IoT | TAAL |
P |
2017 |
Owfuzz(Cao, 2021) | ● | Wi-Fi | Manual |
P | State | ||
2022 |
Meng et al.(Meng et al., 2021) | ◐ | General | Manual |
PG |
P |
State |
2022 |
Greyhound(Garbelini et al., 2022c) | ◐ | Wi-Fi | Manual |
SPHS |
P & S |
State & # of Bugs |
2022 |
SGFuzz(Ba et al., 2022) | ◐ | General | PAL |
SRMS |
P |
State & Code Cov |
2022 |
Braktooth(Garbelini et al., 2022a) | ◐ | Bluetooth | TAAL |
SPHS |
P |
State Transition |
2022 |
L2Fuzz(Park et al., 2022) | ● | Bluetooth | Manual |
Sequential |
P |
State Cov |
2022 |
AmpFuzz(Krupp et al., 2022) | ◐ | UDP | - |
P |
BAF |
2022 |
FUME(Pearson et al., 2022) | ● | MQTT(OASIS, 2019) | - |
P |
Response Freshness |
2022 |
Garbelini et al.(Garbelini et al., 2022b) | ● | 4G/5G | TAPL |
P & S |
2023 |
FeildFuzz(Bytes et al., 2022) | ● | Codesys | - |
P |
Code Cov |
2023 |
BLEEM(Luo et al., 2023) | ● | General | TAAL |
SRHS |
P & S |
State |
2023 |
Tyr(Chen et al., 2023) | ◐ | Blockchain | Manual |
SRHS |
P |
State & Code Cov |
2023 |
CHATAFL(Meng et al., 2024) | ◐ | General | LLM |
LLM |
P |
State & Code Cov |
2023 |
EmNetTest(Amusuo et al., 2023) | ◐ | General | Manual |
Sequential |
P & S |
State |
2023 |
DYFuzzing(Ammann et al., 2024) | ◐ | General | Manual |
SPMS |
P & S |
Code Cov & # of Bugs |
2023 |
FuzzBD(Kim et al., 2023) | ● | USBPD | Manual |
P |
State |
2023 |
Mallory(Meng et al., 2023) | ◐ | DS | TAAL |
SPMS |
P & S |
Event Trace |
○: Whitebox Fuzzer; ●: Blackbox Fuzzer; ◐: Greybox Fuzzer; T: Taxonomy; CMC: Communication Model Construction; General: The fuzzer is not designed for a specific type of protocol; DS: Distributed System; PAL: Program-Analysis-assisted Learning; TAAL: Traffic-Analysis-based Active Learning; TAPL: Traffic-Analysis-based Passive Learning; SRMS: State Rarity-preferred Monolithic Scheduling; SPHS: State Performance-preferred Hierarchical Scheduling; SCHS: State Complexity-preferred Hierarchical Scheduling; SPMS: State Performance-preferred Monolithic Scheduling; SRHS: State Rarity-preferred Hierarchical Scheduling; PG: Property-Guided; -: Not implemented; GUI Ops: GUI Operations; BAF: Bandwidth Amplification Factor; P: Packet level construction; S: Sequence level construction.
@@ 4.1. Communication Model Construction
To empower a traditional fuzzer with semantic constraint knowledge, most existing works model the communication as state machines or their variants to guide fuzzing. A state machine is a data structure that describes the internal state transitions of a protocol implementation. State machines are intrinsically the supergraphs of directed graphs, such as Deterministic Finite Automaton (DFA) or Mealy machine (de Ruiter and Poll, 2015; Somorovsky, 2016; Walz and Sikora, 2020; Fiterau-Brostean et al., 2020; Garbelini et al., 2020; Ruge et al., 2020; Ren et al., 2021; Garbelini et al., 2022c). The nodes and edges in the state machine represent the internal state of the entity and the transitions caused by receiving or sending a message, respectively. By referring to the state machine, protocol fuzzers can be aware of the current target state, and can generate testcases according to the message types that are acceptable in the current state, thereby improving the effectiveness of testcases. Note that a protocol implementation may have multiple communication models, as it may behave differently depending on its working mode or configurations. For example, a Wi-Fi device can be configured to run in AP-mode, STA-mode, or P2P-mode (Xie et al., 2020; Cao, 2021) and a SIP implementation can be configured as a client, server, or proxy (Ozavci, 2013), each of them react differently to requests, thus having a different communication model. Most existing works treat these implementations running in different configurations as different targets: they construct one communication model for one given configuration.
In this section, we provide a taxonomy of the existing works based on their communication model construction methods. As shown in Fig 5, they are divided into two categories: (i) top-down approaches, and (ii) bottom-up approaches.

### v3

@@ 3.1.2. Constrained Testing Environment
Protocol fuzzing usually faces a constrained testing environment due to the tight coupling between protocols and the hardware. Firstly, numerous protocols are designed for communications between low-level physical devices or for communications occurring in specialized sectors, such as protocols residing in the lower layers of the OSI reference model, i.e., the physical and data link layers (Chen et al., 2005; 802, 2021, 2020; Alliance, 2015; SIG, 2016), or protocols designed for specific sectors such as automotive (Zuo et al., 2021; AUTOSAR, 2016; OMG, 2018), industrial control system (ICS) (Zhao et al., 2019; Bytes et al., 2022), electrical grids, and aviation systems. In these cases, the testing throughput will be limited by the hardware dependencies, such as the lack of auotmation (Park et al., 2022; Song et al., 2022), the bottleneck for scalable fuzzing (Bytes et al., 2022; Ruge et al., [n. d.]), etc. In addition, these physical dependencies also limit the application of advanced fuzzing techniques. This is because many advanced fuzzing techniques require greybox or whitebox testing information from the test target, which cannot be satisfied due to the lack of program analysis frameworks on these specific hardware (Zuo et al., 2022; Schepers et al., 2021; Zuo et al., 2021).
@@ 3.2. Summary of Existing Protocol Fuzzers
A general fuzzer consists of three basic components, namely input generator, executor, and bug collector. In one fuzzing iteration, input generator first produces a testing input to executor. Then executor executes the PUT with the given input and collects runtime information for the other two components. Finally, the bug collector checks the runtime information to determine whether the input has triggered a bug.
The unique characteristics of protocols introduced in Section 3.1 present distinct challenges in the design of these three components. The high communication complexity demands that the input generator not only adheres to semantic constraints, including both intra- and inter-message constraints, but also generates inputs from a multitude of testing perspectives to enhance the discovery of vulnerabilities. Similarly, high communication complexity also poses challenges for the bug detector, requiring it to detect whether test cases violate various security properties of the communication process. Furthermore, the constraints of the testing environment impose limitations on the executor’s scalability and its capability to collect runtime information, further influencing the bug collector’s reliance on a limited set of bug detection mechanisms. These challenges underscore the need for specialized adaptations in the fuzzing framework to effectively address the intricacies of protocol testing.
We have analyzed current protocol fuzzing research and encapsulated these efforts into a technical framework for protocol fuzzers, as depicted in Fig. 4. Note that existing protocol fuzzing work still follows the high-level concepts of general fuzzing but proposes specific enhancements in these subcomponents to solve protocol-specific challenges.
Input generator. Ideally, this component is responsible for generating inputs to expose vulnerabilities inside PUTs as effectively as possible. To realize this, protocol fuzzers usually implement input generators with three main subphases including communication model construction, scheduling and testcase construction.
@@
Bug collector. The primary objective of the bug collector component is twofold: to increase the diversity of bug types it can detect and to enhance the accuracy of these detections. The component is finely tuned to meticulously identify a broad spectrum of vulnerabilities, ranging from memory-safety bugs like buffer overflows to more subtle non-memory-safety bugs such as logic errors and specification violations.
@@ 4. Input Generator
In this section, we will introduce in detail how the existing work improves the input generator to solve the unique challenges in protocol fuzzing. As shown in Table LABEL:table-inputgen, we summarize the techniques used in existing work in designing the three key phases in input generator. The covered works are selected from our paper set as long as their main techniques are directly related to the input generator. In addition to the statistics of the mentioned three phases for these works, the table also lists their feedback information. According to Fig. 4, the feedback information can be provided by the executor or the bug collector. We only discuss how the feedback information is used in input generator but leave the details related to the feedback collection for Section 5.2.
| Years | Work | Tax | Target |
Scheduling |
Feedback | ||||
| 2013 | BED(Sesterhenn and Muench, 2013) | ● | General | Manual | Sequential | P | State | ||||
| 2013 | Tsankov et al. (Tsankov et al., 2013) | ● | General | Manual | - | P & S | - | ||||
| 2014 | Peach(Eddington, 2014) | ● | General | Manual | Sequential | P | State | ||||
| 2015 | Pulsar(Gascon et al., 2015) | ● | General | TAPL | SCHS | P | State | ||||
| 2015 | Beurdouche et al.(Beurdouche et al., 2015) | ● | TLS | Manual | Random | S | State | ||||
| 2015 | Ruiter et al.(de Ruiter and Poll, 2015) | ● | TLS | TAAL | Random | S | State | ||||
| 2016 | TLS-Attacker(Somorovsky, 2016) | ● | TLS | Manual | Random | P & S | - | ||||
| 2016 | Driller(Stephens et al., 2016) | ○ | General | - | SPMS | P | Code Cov | ||||
| 2017 | Fan et al.(Fan and Chang, 2018) | ● | General | TAPL | - | P & S | - | ||||
| 2017 | WiFuzz(Vanhoef, 2017) | ● | Wi-Fi | Manual | Sequential | P & S | - | ||||
| 2018 | TCPWN(Jero et al., 2018) | ● | TCP | Manual | Sequential | P & S | State | ||||
| 2018 | IoTFuzzer(Chen et al., 2018) | ● | IoT (OASIS, 2019; Shelby, 2014) | - | - | P | - | ||||
| 2018 | Danial et al.(Daniel et al., 2018) | ● | OpenVPN(openvpn, 2014) | Manual | Random | S | - | ||||
| 2018 | DELTA(Lee et al., 2018) | ● | OpenFlow(Foundation, 2015) | Manual | - | P & S | - | ||||
| 2019 | SeqFuzzer(Zhao et al., 2019) | ● | ICS | TAPL | - | P | State | ||||
| 2019 | Polar(Luo et al., 2019) | ◐ | ICS | - | SPMS | P | Code Cov | ||||
| 2019 | IoTHunter (Yu et al., 2019) | ◐ | IoT | TAAL | Sequential | P | Code Cov | ||||
| 2019 | MoSSOT (Shi et al., 2019) | ● | SSO(Group, 2018) | Manual | Sequential | P & GUI Ops | - | ||||
| 2019 | Chen et al. (Chen et al., 2019) | ◐ | General | - | SPHS | P | State & Code Cov | ||||
| 2019 | Fuzzowski(Romero and Rivas, 2019) | ◐ | General | Manual | - | P | State & Code Cov | ||||
| 2020 | Exploiting Dissent(Walz and Sikora, 2020) | ● | TLS | - | Random | P | - | ||||
| 2020 | DTLS-Fuzzer(Fiterau-Brostean et al., 2020, 2022) | ● | DTLS | TAAL | Random | S | State | ||||
| 2020 | AFLNET (Pham et al., 2020) | ◐ | General | TAAL | SRHS | P | State & Code Cov | ||||
| 2020 | SweynTooth (Garbelini et al., 2020) | ● | BLE(SIG, 2016) | Manual | SPHS | P & S | State & # of Bugs | ||||
| 2020 | Frankenstein (Ruge et al., [n. d.]) | ◐ | Bluetooth | Manual | Random | P & S | State & Code Cov | ||||
| 2020 | Peach* (Luo et al., 2020) | ◐ | ICS | - | - | P | Code Cov | ||||
| 2020 | aBBRate (Peterson et al., 2020) | ● | TCP | Manual | Sequential | S | State | ||||
| 2020 | IJON (Aschermann et al., 2020) | ◐ | General | PAL | Random | P | State & Code Cov | ||||
| 2020 | FuSeBMC (Alshmrany and Cordeiro, 2020) | ○ | General | - | Sequential | P | Code Cov | ||||
| 2020 | DPIFuzz (Reen and Rossow, 2020) | ● | QUIC(Iyengar and Thomson, 2021) | Manual | Random | P & S | - | ||||
| 2020 | Zou et al.(Zou et al., 2020) | ● | General | - | - | P | - | ||||
| 2021 | ICS3Fuzzer (Fang et al., 2021) | ● | ICS (Corporation, 2020; Electric, 2009) | PAL | SCHS | P & GUI Ops | State | ||||
| 2021 | StateAFL (Natella, 2022) | ◐ | General | PAL | SPHS & SRHS | P |
| 2021 | TCP-Fuzz (Zou et al., [n. d.]) | ◐ | TCP | Manual | Random | P & S & Syscall | State Transition | ||||
| 2021 | Snipuzz(Feng et al., 2021) | ● | IoT | - | - | P | - | ||||
| 2021 | Z-Fuzzer(Ren et al., 2021) | ◐ | Zigbee | - | - | P & Interrupt | Code Cov | ||||
| 2021 | PAVFuzz(Zuo et al., 2021) | ◐ | AV (AUTOSAR, 2016; OMG, 2018) | Manual | Sequential | P | Code Cov | ||||
| 2021 | Aichernig et al.(Aichernig et al., 2021) | ● | IoT | TAAL | - | P | - | ||||
| 2017 | Owfuzz(Cao, 2021) | ● | Wi-Fi | Manual | - | P | State | ||||
| 2022 | Meng et al.(Meng et al., 2021) | ◐ | General | Manual | Property-Guided | P | State | ||||
| 2022 | Greyhound(Garbelini et al., 2022c) | ◐ | Wi-Fi | Manual | SPHS | P & S | State & # of Bugs | ||||
| 2022 | SGFuzz(Ba et al., 2022) | ◐ | General | PAL | SRMS | P | State & Code Cov | ||||
| 2022 | Braktooth(Garbelini et al., 2022a) | ◐ | Bluetooth | TAAL | SPHS | P | State Transition | ||||
| 2022 | L2Fuzz(Park et al., 2022) | ● | Bluetooth L2CAP | Manual | Sequential | P | State Cov | ||||
| 2022 | AmpFuzz(Krupp et al., 2022) | ◐ | UDP | - | - | P | BAF | ||||
| 2022 | FUME(Pearson et al., 2022) | ● | MQTT(OASIS, 2019) | - | - | P | Response Freshness | ||||
| 2022 | Garbelini et al.(Garbelini et al., 2022b) | ● | 4G/5G | TAPL | - | P & S | - | ||||
| 2023 | FeildFuzz(Bytes et al., 2022) | ● | Codesys v3(NOCHVAY, 2019) | - | - | P | Code Cov | ||||
| 2023 | BLEEM(Luo et al., 2023) | ● | General | TAAL | SRHS | P & S | State | ||||
| 2023 | Tyr(Chen et al., 2023) | ◐ | Blockchain | Manual | SRHS | P | State & Code Cov | ||||
| 2023 | CHATAFL(Meng et al., 2024) | ◐ | General | LLM | LLM | P | State & Code Cov | ||||
| 2023 | EmNetTest(Amusuo et al., 2023) | ● | General | Manual | Sequential | P & S | State | ||||
| 2023 | DYFuzzing(Ammann et al., 2024) | ◐ | General | Manual | SPMS | P & S | Code Cov & # of Bugs | ||||
| 2023 | FuzzPD(Kim et al., 2023) | ● | USBPD | Manual | - | P | State | ||||
| 2023 | Mallory(Meng et al., 2023) | ◐ | Distributed Sys | TAAL | SPMS | P & S | Event Trace |
○: Whitebox Fuzzer; ●: Blackbox Fuzzer; ◐: Greybox Fuzzer; Tax: Taxonomy; General: The fuzzer is not designed for a specific type of protocol; PAL: Program-Analysis-assisted Learning; TAAL: Traffic-Analysis-based Active Learning; TAPL: Traffic-Analysis-based Passive Learning; SRMS: State Rarity-preferred Monolithic Scheduling; SPHS: State Performance-preferred Hierarchical Scheduling; SCHS: State Complexity-preferred Hierarchical Scheduling; SPMS: State Performance-preferred Monolithic Scheduling; SRHS: State Rarity-preferred Hierarchical Scheduling; -: Not implemented; GUI Ops: GUI Operations; BAF: Bandwidth Amplification Factor; P: Packet level construction; S: Sequence level construction.
@@ 4.1. Communication Model Construction
To enhance traditional fuzzers with semantic constraints, it is essential to develop a communication model of the protocol to guide the fuzzing process. This communication model includes both a state model and a data model. The state model details the transitions between different protocol states, while the data model specifies the format and structure of the messages that drive these state transitions. Typically, existing works represent the protocol’s communication model as a state machine or its variants. A state machine is a data structure that describes the internal state transitions of a protocol implementation. State machines can be represented as directed graphs, such as a deterministic finite automaton (DFA) or a Mealy machine (de Ruiter and Poll, 2015; Somorovsky, 2016; Walz and Sikora, 2020; Fiterau-Brostean et al., 2020; Garbelini et al., 2020; Ruge et al., [n. d.]; Ren et al., 2021; Garbelini et al., 2022c). In these graphs, nodes represent the internal states of an entity, while edges represent state transitions caused by receiving or sending certain types of messages. By referring to the communication model, protocol fuzzers can be aware of the current target state, and can generate testcases according to the data model of message types that are acceptable in the current state, thereby improving the effectiveness of testcases. Note that a protocol implementation may have multiple communication models, as it may behave differently depending on its working mode or configurations. For example, a Wi-Fi device can be configured to run in AP-mode, STA-mode, or P2P-mode (Xie et al., 2020; Cao, 2021) and a SIP implementation can be configured as a client, server, or proxy (Ozavci, 2013), each of them reacts differently to requests, thus having a different communication model. Most existing work treats these implementations running in different configurations as different targets: They construct one communication model for one given configuration.
In this section, we provide a taxonomy of the existing work based on its communication model construction methods. As shown in Fig. 5, they are divided into two categories: (i) top-down approaches, and (ii) bottom-up approaches.